OpenStudy (anonymous):

A family has four siblings, each born exactly 4 years after the other. If the sum of their ages is 128, how old is each person? Define each variable first. Show equation(s) used to solve.

OpenStudy (anonymous):

(n)+(n+4)+(n+8)+(n+12)=128 4n+24=128 4n=104 n=26 The first child is n = 26 The second child is n+4 = 30 The third child is n + 8 = 34 The fourth child is n +12 = 38. To check: 26+30+34+38=128 128=128✓

OpenStudy (anonymous):

awesome! Thank you so much!